Ships A and B leave port together. For the next two hours, ship A travels at 30 mph in a direction 34 ∘ west of north while ship B travels 26 ∘ east of north at 31 mph .

What is the distance between the two ships two hours after they depart?

What is the speed of ship A as seen by ship B?
Answer #1

here,

the velocity of ship A , vA = 30 mph * ( - cos(34) i + sin(34) j)

vA = - 24.9 mph i + 16.78 mph j

the velocity of ship B , vB = 31 mph * ( sin(26) i + cos(26) j)

vB = 13.59 mph i + 27.86 mph j

the speed of ship A relative to ship B , Vab = vA - vB

Vab = - 38.49 mph i - 11.08 mph j

the magnitude of relative velocity , |Vab| = sqrt(38.49^2 + 11.08^2) = 40.05 mph

direction , theta = arctan((-11.08)/(-38.49))

theta = 16.06 degree South of West

the distance between the two ships two hours after they depart, s = |vab| * 2 miles

S = 80.1 miles